Q3 m Please find out the force Frequired to pull the mass out to angle o. (10pts)
Answer #1

here, for the ball to be in equilibrium

balancing the forces in horizontal direction

T * cos(θ) = m * g

balancing the forces in vertical direction

T * sin(θ) = F

>> sin(θ)/cos(θ) = F/(mg)

F = m*g * tan(θ)

hence , the force F is m*g * tan(θ)